Technical Observing: 29 December 2012 UT

Observer: DMiller, JHill (LBTO)
Telescope Operator: GBechetti (LBTO)
Software Support: TSargent (Tucson)
AO Support: DMiller (LBTO)
Telescope Support: JUrban (LBTO)
Instrument Support: ESolheid (LBTO)

Summary

  • Phase 2 (w/o dynamic reconstructor) sextractor GCS v2.7.3 B4 active optics has been tested on-sky with MODS1 and LUCI1 and released for science use. Convergence in the presence of large aberrations is improved, and sensitivity to faint stars with high background is improved.

  • PCS patch PCSNov21_e in 2012H has been certified on-sky for monocular science use (not yet tested in binocular mode)

  • We've gained some ground on Issue 4251 (rows or columns at edge of acq image), but have not completely solved it.

  • General parameters have been set for the 180 deg rotated AGW2, but DX M3 problems have postponed a precise alignment. Only the first steps of that AGW2 alignment work have been completed.

  • Near the end of the night we obtained field aberration measurements with the "new" (20121218) collimation model for LUCI1@LFBG. The result for BS9152 field_5 was that the binodal astigmatism was zero within the measurement errors. These data were taken at an Elevation of 55 degrees. MODS1 @ LDG data was taken in poorer seeing, and was less conclusive. Still to be resolved is the question of whether binodal astigmatism changes as a function of elevation.

  • We tested the higher (medium) order active optics corrections for Z12-Z22 on-sky with both AGW1 and MODS1. The signs have been set, but these have not been deployed for science due to problems with high forces panicing the primary mirror.

Details

Open

23:30 Geno sent test preset, open vent doors and start ventilation. Outside temp=-8C, Chamber temp=1C, Mirror temp=-1C

00:15 Geno opens chamber doors

00:25 John changed RIGHTZEROPOINT for RFBG from -2.4 to 177.6 in PCSInstrument.conf. Restarted PCS

00:35 Doug changed the pointer for GCS configuration files from GCS.2012F to GCS.newWFS. Restarted both GCSs.

Authorized IRTC @ RFBG

00:50 Preset to M5 star. Not found. Spiral search. Nothing

01:05 Nothing found. Geno reset encoders again. He already had reset them earlier. Doug went chamber and checked that tertiary was in position, to within a few degrees.

01:25 No star found with BIG spiral search (5 arcmin diameter). Switch to LFBG.

Authorized LUCI @ LFBG

01:30 Preset and find star on the acquisition image. Geno centers it.

01:35 Preset to BS star to collimate. All WFSC images are blank (zeros). Geno will reset AzCam servers etc while we work on right side.

Authorized IRTC @ RFBG

01:47 Authorize RFBG. Preset to 1 mag star. Will move tertiary selector by 0.1 degree increments to see if we can find star. Enable RZ collimation.

01:50 Move in positive (+) direction. See dim stars (so we are seeing some sky).

01:57 Move in negative (-) direction because John thought there was a gradient in the background at 0 deg, and not at 0.6. Went to -1.4 deg.

02:20 Move in positive direction again up to 2.6. Had Geno home the secondary. Dim star still on acquisition image We continued to +3.4 deg in steps of 0.2 deg.

02:25 Look at pupil position on wfs images

selector position image center of pupil
3.4 1 10,14
4.0 2 9.5,13
4.6 3 9,12
5.8 4 8,10
7.0 5 7,8
     

02:42 Preset to dim star in field. Acquisition moved the star in the correct direction. Guiding kept the star in the same position, but because of drift (star not really on-axis) the guiding could not keep up.

02:47 Send TRACK preset. Focus changed quite a bit. Why?

02:55 RZglobal=5.12 deg RXabsolute=-20.92 deg ----> Found a Bright star.

03:20 John moved M2 Xglobal by -1.2 mm to get off an +RY limit of DX M2

03:35 Coma was diverging, so had to change WFS_rotangleOffs from 180 to 0. Not sure why...

High Order Active Optics Corrections

03:40 Doug is adjusting the signs of Z12-Z21 aberrations on-sky. (in parallel with John's offset tests)

Offset Tests on DX

03:40 offset 0 10 (RADEC ABS) works OK
offset 0 0
offset 60 0
offset 0 0
offset 50 0 (DETXY REL) works OK
offset 0 ??

Conclude that offsetting is working fine with these settings (not yet verified on IRTC).

Preset Tests on DX

Did a series of on- and off-axis presets to verify active optics rotation and transform. It basically works, but we have a huge (>20 pixel) pupil wobble that is inducing high order aberrations. The is because of the very rough alignment of DX M3.

04:40 DX Primary mirror panics. (maybe from the high order aberrations?)

SX Field Aberration Measurements at LFBG

04:45 Changed link for PCS executable from PCS08Nov_d to PCS21Nov_e. Stopped and restarted PCS

04:50 Preset to BS9122. This preset and an offset have verified that this version of PCS is monocular-safe. We'll keep it in even though we haven't done the binocular tests yet.

Data is in /home/aoeng/data/20121229/field_1 for BS9122 at EL=59 deg.

05:29 GCS is having problems with sextractor picking the moonlit edge of the acq field. So we are moving to BS9107 which is away from the moon.

05:30 Applied the Issue 4251 ROI parameters to LUCI_L.cfg.

BS9107 = field_1

Data is in /home/aoeng/data/20121229/field_1 for BS9107 at EL=74 deg.

05:38 Star missed the acq field.

05:40 Send acquire preset to recenter

06:40 Stopped and restarted GCS to absorb the ROI parameters. This removed the overscan region at the top of the image, but didn't fix the columns at the left or the bias ramp at the bottom.

06:56 Set X1 in ROI params from 8 to 12 pixels.

07:03 Set X1 in ROI params from 12 to 9 pixels.

SX Field Aberrations for MODS1 @ LDG

Switch to MODS1 to check things out there.

07:16 woke up MODS agw

Had to reboot azcamserver and azcamg to get an acquisition image.

Updated guide correction factor to 1.0 in LUCI_L.cfg for the new sextractor GCS directory.

BS9122 = field_2

08:19 field_collect on BS9122 with MODS1, but preset fails to find the star, even though we were just observing it moments before.

08:22 ptreset and home M2 hexapod, but we don't understand why there was a big change in pointing.

08:37 field_collect on BS9122 with MODS1

08:40 preset fails to offset star into MODS hotspot. Why

This scenario has +345 arcsec RY dialed in from active optics due to the initial coma. Using the venerable MODS collimation model.

08:43 rerun field_collect

08:45 same error again, so edit the large radius stars out of the field list.

08:49 rerun field_collect EL=71 deg

star 20 is a binary AND too close to the edge of the field.

08:53 rerun field_collect

08:58 finally getting some data points. Seeing overall is worse 0.9-1.4 arcsec on guiding, but dome seeing is much lower than the earlier LUCI field data.

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   8.9", Delta IE =  -8.9"
         Tilt =  46.9", Delta CA = -46.9"
  
                M2 Mis-alignment 
             M2 Arcseconds and Millimeters
         -------------------------------------------
          Delta RX =    88.", Delta Y =  0.455 mm
          Delta RY =  -461.", Delta X = -2.397 mm

Note, the measured RY for MODS may be the opposite sign for LUCI because of the additional flip due to the tertiary.

BS9122 = field_3

Until now (for most of the Fall 2012) MODS1 has been using collimation models from 20111225.

Since 20121223 LUCI has been using SXSM model from 20121218, and SXPM model from 20120927.

Changed MODS1 M1 collimation model from SXPMMODSCollimation.20111225.dat to SXPMLUCIFERCollimation.20120927.dat
Changed MODS1 M2 collimation model from SXSMMODSCollimation.20111225.dat to SXSMLUCIFERCollimation.20121218.dat

09:35 Initialize PSFL to read the new models.

This has some initial coma and focus, but active optics drives RY to +63 rather than +400 arcsec with the previous model.

Now take another set of MODS field data in this configuration. Seeing is still around 1.3 arcsec at EL=61, but dome seeing not too bad.

09:49 field_collect on BS9122 for MODS1 with LUCI's collimation model installed.

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  34.8", Delta IE = -34.8"
         Tilt = -15.3", Delta CA =  15.3"
  
                M2 Mis-alignment 
             M2 Arcseconds and Millimeters
         -------------------------------------------
          Delta RX =   342.", Delta Y =  1.778 mm
          Delta RY =   150.", Delta X =  0.782 mm

Testing Z12-Z22 corrections with MODS1

10:15 Turn on the high order corrections. WFS_highfactor = 0.5

10:25 Well collimated high order (12-22). Could still see structure in the guider image, due to variable seeing.

BS9135 = field_4

10:31 ACQUIRE preset to BS9135

10:35 field_collect SX M2 RX=249 RY=47

10:45 MODS field data at 43 elevation and "new" (20121218) LUCI collimation model

11:10 Shell RIP. Was the last off-axis star compromised?

11:26 restart field_collect on line 6, but 8 min delay to John not paying attention to pop-up window.

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   7.1", Delta IE =  -7.1"
         Tilt =  21.2", Delta CA = -21.2"
  
                M2 Mis-alignment 
             M2 Arcseconds and Millimeters
         -------------------------------------------
          Delta RX =    69.", Delta Y =  0.361 mm
          Delta RY =  -208.", Delta X = -1.080 mm

11:30 Switch back to RFBG

Authorize LUCI @ LFBG

BS9152 = field_5

12:00 Preset to BS9152

12:05 field_collect for LUCI field data at 55 elevation (limit=300)

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  -5.5", Delta IE =   5.5"
         Tilt =  -8.0", Delta CA =   8.0"
  
                M2 Mis-alignment 
             M2 Arcseconds and Millimeters
         -------------------------------------------
          Delta RX =   -54.", Delta Y = -0.278 mm
          Delta RY =    79.", Delta X =  0.411 mm

BS9152 = field_6

12:40 LUCI field data at 60 elevation. Added Tip=0 Tilt=-40 on SX M2, CA=+40, RY=+455 from +63 arcsec.

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -10.5", Delta IE =  10.5"
         Tilt = -49.6", Delta CA =  49.6"
  
                M2 Mis-alignment 
             M2 Arcseconds and Millimeters
         -------------------------------------------
          Delta RX =  -104.", Delta Y = -0.539 mm
          Delta RY =   487.", Delta X =  2.533 mm

BS9152 = field_7

13:03 LUCI field data at 65 elevation. Added Tip=-40 Tilt=0 on SX M2, IE=+40, RY=-22 from 325. (had restored Tilt and CA to original values)

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -12.8", Delta IE =  12.8"
         Tilt = -23.6", Delta CA =  23.6"
  
                M2 Mis-alignment 
             M2 Arcseconds and Millimeters
         -------------------------------------------
          Delta RX =  -126.", Delta Y = -0.655 mm
          Delta RY =   232.", Delta X =  1.204 mm

BS9152 = field_8

13:28 LUCI field data at 66 deg elevation. Tip=0 Tilt=0 The parameters are the same as set field_5 at 55 elevation. (had restored Tip and IE to original values)

13:46 End of data collection for impending dawn. No stars could be seen in the wfs image.

Note only thee points, and it was getting bright outside.

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  16.2", Delta IE = -16.2"
         Tilt = -18.1", Delta CA =  18.1"
  
                M2 Mis-alignment 
             M2 Arcseconds and Millimeters
         -------------------------------------------
          Delta RX =   159.", Delta Y =  0.826 mm
          Delta RY =   178.", Delta X =  0.924 mm

Close

-- DougMiller - 29 Dec 2012
Topic revision: r9 - 11 Jan 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback